OpenStudy (anonymous):

If f and g are functions of time, and at time t = 3, f equals 9 and is rising at a rate of 6 units per second, and g equals 7 and is rising at a rate of 9 units per second, then f/g equals_____and is changing at a rate of______units per second.

OpenStudy (anonymous):

Use the quotient rule \[ \left ( \frac f g\right)^' = \frac { f' g - g'f} {g^2} \]

OpenStudy (anonymous):

erm what would the equation look like if i were to use the quotient rule

OpenStudy (anonymous):

would it be 9^6/7^9

OpenStudy (anonymous):

or 9x^6/7x^9

OpenStudy (anonymous):

You are givien f(3), f'(3), g(3) and g'(3). Replace in the formula.